Two forces, F1 = 8.0 N, 60o Northwest and F2 = 12.0 N, Southwest, are applied to a particle at the origin. What third force is needed so that the particle will not be displaced from the origin?


1
Expert's answer
2021-11-22T16:20:16-0500

"F_x=-8\cdot\cos60°-12\cdot\cos45°=-12.5\ (N)"


"F_y=8\cdot\sin60°-12\cdot\sin45°=-1.6\ (N)"


"F=\sqrt{12.5^2+1.6^2}=12.6\ (N)"


"\tan\theta=\frac{1.6}{12.5}=0.128\to \theta=7.3°"N of E.